A company has current assets of $90,000 and its current ratio is 2.5. Assume that the company pays off some accounts payable in the amount of $4,000. The current ratio after this transaction is closest to:

                  A. 2.39

                  B. 2.69

                  C. 2.50

                  D. 2.81
Answer #1

Current ratio=current assets/Current liabilities

Hence Current liabilities now=(90,000/2.5)=$36,000

Journal would be:

Accounts payable a/c...Dr $4000

To Cash $4000

Hence Current liabilities would reduce and would become=(36000-4000)=$32000

Also current assets would be=(90,000-4000)=$86,000

Hence current ratio would be=86,000/32,000

=2.69(Approx).